An aircraft carrier left Diego Garcia and traveled toward St. Vincent at an average speed of 12 km/h. A container ship left one hour later and traveled in the same direction but with an average speed of 15 km/h. Find the number of hours the aircraft carrier traveled before the container ship caught up.
Answer by mananth(16949)

An aircraft carrier left Diego Garcia and traveled toward St. Vincent at an average speed of 12 km/h.
Let the time taken to meet be x
distance travelled = 12x

A container ship left one hour later and traveled in the same direction but with an average speed of 15 km/h.
one hour later
time = (x-1) hours
distance travelled = 15(x-1)
distance travelled by both when they meet is same
15(x-1) = 12x
15x-12x = 15
3x= 15
x= 5